A random sample of 100 recorded deaths in the country in 2011 showed an average lifespan of 71.8 years with a standard deviation of 8.9 years. Does this seem to indicate that the average lifespan today is greater than 70 years? Use 0.05 as level of significance.


1
Expert's answer
2021-12-07T09:55:19-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 70"

"H_1:\mu>70"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

"df=n-1=100-1=99" degrees of freedom,and the critical value for a right-tailed test is "t_c = 1.660391."

The rejection region for this right-tailed test is "R = \{t: t > 1.660391\}."

The t-statistic is computed as follows:


"t=\dfrac{x-\mu}{s/\sqrt{n}}=\dfrac{71.8-70}{8.9/\sqrt{100}}\approx2.022472"

Since it is observed that "t = 2.022472 >1.660391= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for "df=99" degrees of freedom, "t=2.022472", right-tailed from Student T-Value Calculator is the Significance Level "p = 0.022912," and since "p = 0.022912<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "70," at the "\alpha = 0.05" significance level.